# Agent-Based Insight into Eco-Choices: Simulating the Fast Fashion Shift

### v1

@@ Abstract
Fast fashion, Agent-Based Modeling, Demand, Consumer behavior, Environment, Working conditions.
I Introduction
@@
Consumption of goods and services serves as a medium for expressing and reinforcing our identity to both ourselves and others. Clothing, in particular, is not just mere utility but a visual language that allows individuals to signal their belonging to certain social groups, subcultures, or communities. This phenomenon reaches its apogee in cultures where consumption became the main way of self-actualization and self-expression. The imperative to continually evolve one’s identity through new clothing and the allure of staying fashionable sometimes overshadows concerns about environmental impact, labor conditions, and ethical sourcing. How does this widespread shift in our consumption patterns impact both our world and its inhabitants?
@@
The rise of the fast fashion business model was a response to the growing demand for inexpensive yet trendy clothing. This rapid shift towards lower prices and quality, however, came at the expense of reducing wages, extending the working hours of garment workers, increasing the production of cheaper textiles, and more shortcuts. As a result of these changes in manufacturing practices, the number of times a garment is worn on average has steadily declined, with people accumulating more clothes than ever before.
@@
The prosperity of fast fashion retailers largely depends on consumer behavior, driven by the desire for increased clothing ownership made possible by continually declining prices. Although individual actions are often perceived as insignificant compared to the environmental impact of large corporations, studies have highlighted the importance of the consumer’s role in European policies [26]. This study aims to investigate the impact of awareness and education about sustainability on individuals’ decision-making processes regarding fast fashion purchases. Additionally, we employ Agent-Based Modeling to assess how various influences can alter decision-making and steer public opinion towards more sustainable garment choices. Recognizing that public willingness to endorse and actively participate in initiatives addressing issues like fast fashion is pivotal for progress ([16]), we seek to examine shifts in public opinion through diverse influences such as social media, peer pressure, and governmental intervention.
@@
The rest of this thesis will follow this structure: in Section II, we review existing literature on fast fashion and justify the necessity of this research. Section III briefly explored the date we used. In Section IV, we provide a detailed description of the model, including the methods employed, an explanation of agent-based modeling, and an overview of my model’s structure. Next, Section VI presents an analysis of models and simulations conducted. Finally, in Section VII, we summarize the findings, address limitations, and propose avenues for future research.
II Literature overview
@@
Fast fashion denotes the rapid manufacture of inexpensive, substandard clothing often imitating popular styles from fashion labels, renowned brands, and independent designers [22]. It has risen to prominence in the fashion industry by selling large volumes of apparel at affordable rates, triggering an unprecedented level of garment consumption. In this section, we will examine fast fashion and its adverse effects on the environment, as well as its impact on the individuals involved in production. We will also explore the cultural dynamics that foster the growth of fast fashion, including consumer behavior.
II-A Fashion’s environmental impact
@@
Manufacturing locations and Supply Chain. The relocation of fast fashion manufacturing to countries with higher carbon footprints, such as India and those in Southeast Asia, is primarily motivated by cost-cutting measures [31]. In these countries, emissions control and regulations are often lax, enabling companies to prioritize profit extraction even at significant environmental expense. However, this shift presents significant challenges to energy supply security and compliance with emissions reduction commitments [32]. Often, garment production involves multiple countries, leading to increased logistical steps between processes driven by economic considerations [3]. Additionally, there is a growing reliance on air cargo for transportation, significantly amplifying environmental impact.
@@
II-B Fast fashion and working conditions
@@
On a positive note, the fast fashion sector has contributed to overseas income, foreign exchange earnings, women’s empowerment, and the overall export value and GDP in countries such as Bangladesh [5] and Ethiopia [23]. However, the notion of economic development transcends merely raising the per capita income of industry workers. Coupled with increasing income inequality, the working class often remains trapped in poverty, facing job insecurity, meager wages, extended work hours, limited access to public services, substandard health care, and dismal living and working conditions [24]. They also endure a lack of legal rights and face physical and mental threats to their economic and social well-being.
@@
II-C Throw away culture
The evolution of the fashion business model has been closely related to shifts in consumer attitudes toward clothing. The frequent updates of styles and affordability have fostered a culture of impulse buying and a perception of clothing as easily disposable. Fast fashion brands now introduce up to 52 seasons per year [6], compared to the traditional four seasons, capitalizing on limited collections that quickly go out of style and the planned obsolescence of their products due to their poor quality.
Additionally, several barriers hinder the adoption of more sustainable shopping habits. Sustainable clothing is often perceived as more expensive than fast fashion items, and there persists a belief that sustainably produced clothing lacks attractiveness and is thus considered unfashionable. Social norms also play a significant role in shaping our shopping behavior, with normative expectations influencing how we perceive our peers’ expectations of us. Consequently, even when individuals express concerns about the environmental and social impacts of fast fashion, these concerns do not necessarily imply changes in their shopping habits. This phenomenon, known as the “Ethical Purchasing Gap” [7], suggests that in fashion purchases, ethical considerations may have some influence, but factors like color and style are likely to be more significant determinants.
III Data
@@
IV Methods
IV-A Purpose
We aim to study garments consumption patterns, identify barriers to adopting more sustainable shopping practices, assess responses to educational interventions, and evaluate the impact of policies and external influences. To do so, we employ Agent-Based Modeling (ABM), which is a computational modeling technique utilized to simulate and analyze the behavior and interactions of diverse individual agents at a micro-level, and the resultant outcomes at the macro-level of a system [35].
The model is implemented in the multi-agent programmable environment NetLogo 6.3.0. A copy of the model with supplementary information can be downloaded from the model library of the CoMSES Net website (www.comses.net).
@@ IV-A Purpose
IV-B Decision-making model
@@
From the dataset, we identify several predictor variables and one response variable, which are described in detail in Appendix A-A.
@@
$A_{p}$ $=b_{0}+b_{1}\cdot A_{sex}+b_{2}\cdot A_{age}+b_{3}\cdot A_{env}+b_{4}\cdot A_{exp}$
$\quad+b_{5}\cdot A_{wca}+b_{6}\cdot A_{know}+b_{7}\cdot A_{trust}+b_{8}\cdot A_{access}$
$\quad+b_{9}\cdot A_{freq},$ (1)
where $A_{i}\in[0,1]$ are agents’ attributes and their corresponding definitions can be found in Table I. The coefficients come from the linear regression analysis and can be found in Table II. Variables not considered in this analysis, detailed in Appendix A-A, were discarded due to a lack of any correlation between changes in their values and the response variable. These include: level of education, perceived consumer effectiveness, income.
@@
$b_{3}$0.0179 Environmental concerns
$b_{4}$0.0488 Normative expectations
$b_{5}$0.1783 Working conditions awareness
$b_{6}$0.1414 Education sustainable fashion
@@
TABLE II: Coefficients of the linear regression
IV-C Initializing the agentset
Agents are initialized based on the information from the dataset used for the linear regression. It is used to create 1050 agents. The attributes each agent is assigned are seen in Table I and are those used in the decision-making model. Aside from these, each agent is randomly assigned three additional static variables that represent their levels of sustainability. Each of these are within the range of $[0.1,0.9]$, where 0.1 indicates low susceptibility and 0.9 - high susceptibility
@@ IV-C Initializing the agentset
    Peer influence susceptibility, $S_{pp}$. This variable determines how easily an agent’s opinion is influenced by other agents in its neighborhood. A higher susceptibility value indicates that the agent is less influential on others. Conversely, if an agent has significant influence over its peers, its susceptibility to their influence will be lower [2].
@@
IV-D Agent-Based Model
@@
IV-D1 Peer influence. Non-polarized agentset
The first kind of agentset is titles “non-polarized” and corresponds to a set of individuals whose opinion on topics tends to homogenize over time because these topics are not perceived as controversial or polarizing ones [29]. In both non polarized and polarized agentset, agents are programmed to interact with a subset of the agents in their neighborhood during each time step. Every agent is set to have 5 close friends, which forms their inner circle, and 10 acquaintances, which forms their outer circle. Although in real-life individuals tend to have more connections [12], [20], we base our choice on the size of the agentset to avoid overconnectedness. The close friends are all connected, while the acquaintances are randomly selected and do not necessarily connect to each other. Hence, each agent has a total of 15 connections, and we allow agents to interact with 10 other agents daily. To introduce some variability, we allow agents to interact with $10\pm a$ friends, where $a\in[1,4]$. In this model, we assume that an agent is equally likely to interact with their close friends as with their acquaintances. Additionally, we assume that each interaction with a friend, whether best or close, has an equal effect on the agent’s perception of their opinion.
@@
$A_{t,i}$ $=(1-S_{pp}(A_{i}))\cdot A_{t-1,i}+\frac{S_{pp}(A_{i})}{\sum\limits_{A_{j}=1}^{A_{j}\in peers(A_{i})}1}\cdot$
$\quad\sum\limits_{A_{j}=1}^{A_{j}\in peers(A_{i})}\left[\frac{1}{3}A_{t-1,j}+\frac{2}{3}B_{t-1,j}\right],$ (2)
@@
where $A_{t,i}$ stands for agent’s updated opinion on a topic, $S_{pp}(A_{i})$ represents agent’s susceptibility towards others, $A_{t-1,i}$ represents their opinion at the previous time step, and $A_{t-1,j}$ and $B_{t-1,j}$ represent peers’ opinion and behavior at a previous time step, respectively. Each agent can only be influenced by those in its neighborhood whose susceptibility is lower than that of the agent. This represent the phenomenon seen in real-life interactions, where lower susceptibility tends to correlate with higher influentiality, and vice versa [2]. This variable also allows us to incorporate another important social phenomenon: inexorable individuals, which are not easily persuaded, moved, or affected by normative expectations. Additionally, both peers’ behavior and attitudes are included in this equation ($A_{t-1,j}$ and $B_{t-1,j}$, respectively). This decision was based on the observation that individuals tend to be influenced twice as much by conformity with peers’ behavior as by normative expectations [42]. In other words, what one’s peers do matters more than one’s expected behavior.
IV-D2 Peer influence. Polarized agentset
@@
The second kind of agentset we explored is titled “polarized.” We want to explore peer influence in a scenario where the population is polarized and does not homogenize their views over time, but instead clusters around multiple opinions. We know that consensus is rarely achieved in real-world scenarios regarding polarized topics because individuals’ initial internal opinions, unlike their expressed opinions, remain unchanged by social interaction [11]. In certain societies and countries, fast fashion and its relation to the environment is a polarized topic, and we want to account for such development.
@@
The equation used for social interactions in a polarized society will be similar to that for a non-polarized society. However, the difference lies in the introduction of a tolerance threshold (denoted by $\tau\in[0.05,0.50]$), which represents how tolerant agents are toward opinions that deviate from theirs by more or less than the threshold. This is explained in more detail in the Appendix A-B. In summary, a lower threshold indicates less tolerance towards differing opinions. When an agent encounters an opinion that exceeds their tolerance threshold, the interaction becomes polarizing rather than unifying.
IV-D3 Social media influence.
@@
In real life, social media platforms adapt to each user, tailoring content based on their preexisting preferences and inclinations. Content one sees is content they already consume or are likely to consume. In our model, we implement such mechanisms by incorporating a feedback loop between the agent’s opinion and their social media platform. It is recognized that social media tends to increase polarization by creating “echo chambers” that limit exposure to information contradicting preexisting beliefs [4], and this is particularly the case for politicized topics [27]. Moreover, social media can be biased towards pro-fast fashion or anti-fast fashion, which is something we consider in this model. Our choice to add a bias term rests on the fact that social media is found to both promote consumption [15], [34], but can also be a powerful tool that can be used to promote sustainable habits [33], [41], [47]. As a result, we derive a function that takes an agent’s current opinion as an input an returns an “influence” opinion to the agent. This feedback loop is given by Eq. (3).
$SM\left(A_{t-1}\right)$ $=b\cdot\left(A_{t-1}\right)^{3}+\frac{-3\cdot b}{2}\left(A_{t-1}\right)^{2}$
$\quad+\frac{3\cdot b}{4}\left(A_{t-1}\right)+\left(\frac{1}{2}-\frac{b}{8}\right)+\beta,$ (3)
were $b=50\cdot S_{sm}$, $S_{sm}$ stands for agent’s susceptibility towards social media, and $\beta\in[-0.30,0.30]$ is the bias. The rationale behind the choice of function for this feedback loop is detailed in Appendix A-B. Although the function is uniform for all agents, the output varies for each agent’s opinion, reflecting the individual calibration of social media platforms. Once we have obtained the social media influence level from the function above, we can find the agent’s opinion after using social media, which is given by:
@@
where $A_{t-1}$ represents agent’s opinion at the previous time step, $A_{t}$ is the updated opinion, $S_{sm}$ stands for social media susceptibility of the agent, and $SM$ is the opinion that social media obtained from the feedback loop. The rationale behind this choice of function is detailed in the Appendix A-B.
IV-D4 Government interventions.
@@
$A_{t}$ $=(1-S_{gov})\cdot g(A_{t-1},GOV)\cdot A_{t-1}$
$\quad+S_{gov}\cdot f(A_{t-1},GOV)\cdot GOV,$ (6)
where $GOV$ is the promoted opinion calculated with Eq. (5), $S_{gov}$ is agent’s susceptibility to government’s interventions, and $A_{t-1}$ is the opinion of an agent at time $t-1$. The functions $g$ and $f$ are exactly the same as described above for social media, and Eq. (6) is identical to Eq. (4) explained in Appendix A-B. Similar to social media, we aimed to prevent drastic changes in agents’ opinions and thus chose a more sophisticated function that allowed for a slower evolution of their viewpoints. Additionally, we implement a function that aims to mimic the diminishing impact of repeated exposure, a phenomenon known as “campaign fatigue” [40]. This will affect agents’ susceptibility to government, $S_{gov}$, in a manner represented by Eq. (7).
@@
IV-E Structure overview and scheduling
Once agents and their friend groups are initialized, the simulation begins. It is set to run for 500 time steps. Each time step, agents deterministically interact with some of their neighbors and social media once, and stochastically interact with government campaigns. More information on the simulation can be found in Appendix A-C.
V Model settings

### v2

@@ Abstract
Keywords— F ast fashion, Agent-Based Modeling, Demand, Consumer behavior, Environment, Working conditions.
1 Introduction
@@
Consumption of goods and services functions as a medium for expressing and reinforcing our identity to both ourselves and others. Clothing in particular is not just mere utility but a visual language that allows individuals to signal their belonging to certain social groups, subcultures, or communities. This phenomenon reaches its apogee in cultures where consumption became the main way of self-actualization and self-expression. The imperative to continually evolve one’s identity through new clothing and the allure of staying fashionable sometimes overshadows concerns about environmental impact, labor conditions, and ethical sourcing. How does this widespread shift in our consumption patterns impact both our world and its inhabitants?
@@
The rise of the fast fashion business model was a response to the growing demand for inexpensive yet trendy clothing. This rapid shift towards lower prices and quality, however, came at the expense of reducing wages, extending the working hours of garment workers, increasing the production of cheaper textiles, and other shortcuts. As a result of these changes in manufacturing practices, the number of times a garment is worn on average has steadily declined, with people accumulating more clothes than ever before.
@@
Fast fashion retailers’ success largely hinges on consumer behavior, fueled by the desire to own more clothing as prices steadily decrease. Although individual actions are often perceived as insignificant compared to the environmental impact of large corporations, studies have highlighted the importance of the consumer’s role in European policies ([36]). This study aims to investigate the impact of awareness and education about sustainability on individuals’ decision-making processes regarding fast fashion purchases. Additionally, we employ Agent-Based Modeling to assess how various influences can alter decision-making and steer public opinion towards more sustainable garment choices. Recognizing that public willingness to endorse and actively participate in initiatives addressing issues like fast fashion is pivotal for progress ([22]), we seek to examine shifts in public opinion through diverse influences such as social media, peer pressure, and governmental intervention.
We model two kinds of agentsets: non-polarized, whose opinions tend to homogenize over time, consistent with classical opinion dynamics models ([10], [14]), and polarized, where agents tend to adhere to their initial beliefs and grow stronger in them. In this case, polarization refers to the processes of driving individual behavior apart. We will come back in detail to the subject of polarization, including a literature review, in Section 4.4.2 below. We find that increased communication generally results in positive changes in concerns, with medium communication levels benefiting non-polarized agents and high communication levels favoring polarized agents. For polarized populations, extreme tolerance levels are less effective, with moderate tolerance yielding the best results. Social media significantly affects non-polarized agents, where moderate pro-sustainability exposure enhances positive changes in concerns, and extreme biases can be detrimental. Conversely, polarized agents show less sensitivity to social media, with lower engagement levels yielding better results. Our analysis of government interventions reveals that while both non-polarized and polarized agentsets respond to pro-sustainability campaigns, the effects are more pronounced in non-polarized agents. Polarized agents show greater resistance to anti-sustainability campaigns and exhibit reduced variability in outcomes. Overall, our simulations suggest that while social media and government interventions shape public concerns and opinions, their effectiveness is significantly influenced by the initial polarization levels and communication dynamics within the population.
@@
The rest of this paper will follow this structure: in Section 2, we review existing literature on fast fashion and justify the necessity of this research. Section 3 briefly explores the data we used. In Section 4, we provide a detailed description of the model, including the methods employed, an explanation of agent-based modeling, and an overview of the model’s structure. Next, Section 6 presents an analysis of models and simulations conducted. Finally, in Section 7, we summarize the findings, address limitations, and propose avenues for future research.
2 Literature overview
@@
Fast fashion denotes the rapid manufacture of inexpensive, substandard clothing often imitating popular styles from fashion labels, renowned brands, and independent designers ([31]). It has risen to prominence in the fashion industry by selling large volumes of apparel at affordable rates, triggering an unprecedented level of garment consumption. In this section, we will examine fast fashion and its adverse effects on the environment, as well as its impact on the individuals involved in production. We will also explore the cultural dynamics that foster the growth of fast fashion, including consumer behavior.
2.1 Fashion’s environmental impact
@@
Manufacturing locations and Supply Chain. The relocation of fast fashion manufacturing to countries with higher carbon footprints, such as India and those in Southeast Asia, is primarily motivated by cost-cutting measures ([43]). In these countries, emissions control and regulations are often lax, enabling companies to prioritize profit extraction even at significant environmental expense. However, this shift presents significant challenges to energy supply security and compliance with emissions reduction commitments ([44]). Often, garment production involves multiple countries, leading to increased logistical steps between processes driven by economic considerations ([2]). Additionally, there is a growing reliance on air cargo for transportation, significantly amplifying environmental impact.
@@
2.2 Fast fashion and working conditions
@@
On a positive note, the fast fashion sector has contributed to overseas income, foreign exchange earnings, women’s empowerment, and the overall export value and GDP in countries such as Bangladesh ([6]) and Ethiopia ([32]). However, the notion of economic development transcends merely raising the per capita income of industry workers. Coupled with increasing income inequality, the working class often remains trapped in poverty, facing job insecurity, meager wages, extended work hours, limited access to public services, substandard health care, and dismal living and working conditions ([33]). They also endure a lack of legal rights and face physical and mental threats to their economic and social well-being.
@@
2.3 Throw away culture
The evolution of the fashion business model has been closely related to shifts in consumer attitudes toward clothing. The frequent updates of styles and affordability have fostered a culture of impulse buying and a perception of clothing as easily disposable. Fast fashion brands now introduce up to 52 seasons per year ([8]), compared to the traditional four seasons, capitalizing on limited collections that quickly go out of style and the planned obsolescence of their products due to their poor quality.
Additionally, several barriers hinder the adoption of more sustainable shopping habits. Sustainable clothing is often perceived as more expensive than fast fashion items, and there persists a belief that sustainably produced clothing lacks attractiveness and is thus considered unfashionable. Social norms also play a significant role in shaping our shopping behavior, with normative expectations influencing how we perceive our peers’ expectations of us. Consequently, even when individuals express concerns about the environmental and social impacts of fast fashion, these concerns do not necessarily imply changes in their shopping habits. This phenomenon, known as the “Ethical Purchasing Gap” ([9]), suggests that in fashion purchases, ethical considerations may have some influence, but factors like color and style are likely to be more significant determinants.
3 Data
@@
4 Methods
4.1 Purpose
We aim to study garments consumption patterns, identify barriers to adopting more sustainable shopping practices, assess responses to educational interventions, and evaluate the impact of policies and external influences. To do so, we employ Agent-Based Modeling (ABM), which is a computational modeling technique utilized to simulate and analyze the behavior and interactions of diverse individual agents at a micro-level, and the resultant outcomes at the macro-level of a system ([47]).
The model is implemented in the multi-agent programmable environment NetLogo 6.3.0. A copy of the model with supplementary information can be downloaded from the model library of the CoMSES Net website ([53]).
@@ 4.1 Purpose
4.2 Decision-making model
@@
From the dataset, we identify several predictor variables and one response variable, which are described in detail in Appendix A.1.
@@
$A_{p}=b_{0}+b_{1}\cdot A_{sex}+b_{2}\cdot A_{age}+b_{3}\cdot A_{env}+b_{4}\cdot A_{exp}+b_{5}\cdot A_{wca}+b_{6}\cdot A_{know}+b_{7}\cdot A_{trust}+b_{8}\cdot A_{access}+b_{9}\cdot A_{freq},$ (1)
where $A_{i}\in[0,1]$ are agents’ attributes and their corresponding definitions can be found in Table 1. The coefficients come from the linear regression analysis and can be found in Table 2. Variables not considered in this analysis, detailed in Appendix A.1, were discarded due to a lack of significant correlation between changes in their values and the response variable. These include: level of education, perceived consumer effectiveness, income. Moreover, while we do include normative expectations and attitude, we do not include behavioral control as suggested by the Theory of Planned Behavior ([1]). This choice is due to two reasons. Firstly, one of most common limitation brought up against buying sustainable clothing is its high cost. However, as further explained in Appendix A.1 and shown in Table (2), neither income nor lack of access to sustainable cloths were significantly correlated with being a sustainable shopper. Secondly, the behavior we aim to affect is not purchasing sustainable clothing, but rather not purchasing clothing from fast fashion brands. This can be seen as buying sustainable clothing, shopping second hand, or simple reducing how much an individual shops for clothing. This is consistent with our definition of the outcome variable.
@@
$b_{3}$ -0.0179 Environmental concerns
$b_{4}$ -0.0488 Normative expectations
$b_{5}$ -0.1783 Working conditions awareness
$b_{6}$ -0.1414 Education sustainable fashion
@@
Table 2: Coefficients of the linear regression
4.3 Initializing the agentset
Agents are initialized based on the information from the dataset used for the linear regression. It is used to create 1050 agents. Once the agentset has been initialized, we create cliques (close friends) and distant links (acquaintances) according to the mechanism described below in Section 4.4.1. The attributes each agent is assigned are seen in Table 1 and are those used in the decision-making model. Aside from these, each agent is randomly assigned three additional static variables that represent their levels of susceptibility. Each of these are within the range of $[0.1,0.9]$, where 0.1 indicates low susceptibility and 0.9 - high susceptibility
@@ 4.3 Initializing the agentset
    Peer influence susceptibility, $S_{pp}$. This variable determines how easily an agent’s opinion is influenced by other agents in its neighborhood. A higher susceptibility value indicates that the agent is less influential on others. Conversely, if an agent has significant influence over its peers, its susceptibility to their influence will be lower. This represent the phenomenon seen in real-life interactions, where lower susceptibility tends to correlate with higher influentiality, and vice versa ([3]).
@@
4.4 Agent-Based Model
@@
4.4.1 Peer influence. Non-polarized agentset
The first kind of agentset is titled “non-polarized” and corresponds to a set of individuals whose opinion on topics tends to homogenize over time because these topics are not perceived as controversial or polarizing ones ([39]). In both non-polarized and polarized agentset, agents are programmed to interact with a subset of the agents in their neighborhood during each time step. Every agent is set to have 5 close friends, which forms their inner circle, and 10 acquaintances, which forms their outer circle. Although in real-life individuals tend to have more connections ([17], [28]), we base our choice of number of connections on the limited size of unique agents (1050) to avoid overconnectedness. The inner circle, comprised of agent’s close friends, is a fully interconnected clique, while the acquaintances are of an agent are not necessarily connected to each other.
Our arrangement can be viewed as a random network with embedded cliques, that builds dense local groups and adds random weak ties to other agents. The structure of our model shares the small-world properties of the Watts–Strogatz network ([62]) — high clustering and short average path lengths. We explicitly embed agents in fully connected cliques and link them to randomly selected others, which allows us to capture the realistic duality of social ties: dense, cohesive friend groups alongside sparse, long-range connections. Unlike Watts–Strogatz, which imposes uniform local connections, or Kleinberg’s model based on spatial proximity ([34]), our network explicitly reflects community structure, making it especially suitable for modeling social behavior within clustered populations. Fashion choices often reflect a clustered population dynamic, where individuals are significantly influenced by their immediate social groups or “cliques.” The patterns introduced by out network structure reflect how individuals frequently look to their immediate social circles for cues on fashion choices, leading to the formation of tightly knit groups with similar fashion preferences.
Hence, each agent has a total of 15 connections. Agents interact with their network at every time step. To introduce variability, we allow agents to interact with $10\pm a$ friends, where $a\sim U[1,4]$. In this model, we assume that an agent is equally likely to interact with their close friends as with their acquaintances. Additionally, we assume that each interaction, whether it is with a close friends or an acquaintance, has an equal effect on the agent’s perception of their opinion. While people may value close friends’ opinions in real life on some topics, it’s unclear if this holds in fashion dynamics. Research suggests individuals tend to follow more influential figures, not necessarily close friends ([21], due to the expert effect ([42]). We account for influential agents with the attribute “Peer influence susceptibility,” depicted in Section 4.3. Moreover, we avoid giving more weight to close friends’ opinions to prevent excessive clustering, which is already introduced by the presence of cliques. Similarly, we do not give more weight to interactions with agents whose opinions align in the non-polarized setting. This differs from previous research that combined agent-based modeling and environmental concerns ([58]), where influence of individuals on each other is a function of their similarity in environmental identity. However, we do introduce an analogous mechanism in polarized agentsets, described in Section 4.4.2.
@@
$A_{t,i}=(1-S_{pp}(A_{i}))\cdot A_{t-1,i}+\frac{S_{pp}(A_{i})}{|C|}\cdot\sum\limits_{j=1}^{|C|}\left[\frac{1}{3}A_{t-1,j}+\frac{2}{3}B_{t-1,j}\right],$ (2)
@@
where $A_{t,i}$ stands for agent’s $i^{\text{th}}$ updated opinion on a topic at time $t$, $S_{pp}(A_{i})$ represents agent’s susceptibility towards others, $A_{t-1,i}$ represents their opinion at the previous time step, $C=\{A_{j}:A_{j}\in peers(A_{i})\}$ is the set of peers the agent interacted with at previous time step $t$, and $A_{t-1,j}$ and $B_{t-1,j}$ represent peers’ opinion and behavior at a previous time step, respectively. It is important to note that an agent can only be influenced by those in its neighborhood whose susceptibility is lower than that of the agent. In other words, an agent is only influenced by more influential agents that the agent themselves is ([4]). On the one had, this allows us to introduce the concept of assimilation: individuals tend to adopt the culture of the dominant agents. The concept of assimilation is further introduced with the state’s influence variable, which is described below. On the other hand, this allows us to incorporate an important social phenomenon: inexorable individuals, which are not easily persuaded, moved, or affected by normative expectations. Additionally, both peers’ behavior and attitudes are included in this equation ($A_{t-1,j}$ and $B_{t-1,j}$, respectively). This decision was based on the observation that individuals tend to be influenced twice as much by conformity with peers’ behavior as by normative expectations ([59]). In other words, what one’s peers do matters more than one’s expected behavior. This choice of opinion updating mechanism is based on the foundational model of opinion dynamics introduced by Morris H. DeGroot in his 1974 paper Reaching a consensus ([14]), where a group of individuals repeatedly update their beliefs by averaging the beliefs of others in their social network. In particular, it closely resembles another well-knows framework in opinion dynamics, the Friedkin–Johnsen (FJ) model, which is an extension of DeGroot’s model. The FJ model allows agents to retain some of their initial opinions.
4.4.2 Peer influence. Polarized agentset
@@
The second kind of agentset we explored is titled “polarized.” Our goal is to explore peer influence in a scenario where the population is polarized, which mean that it does not homogenize their views over time, but instead clusters around multiple opinions. Often times, consensus cannot be achieved in real-world scenarios regarding polarized topics for a variety of reasons. For instance, individuals’ initial internal opinions, unlike their expressed opinions, can remain unchanged by social interaction ([16]). Moreover, polarization seems to burst especially in public discussions evolving around politically and ethically controversial issues ([11], [15]). In certain societies and countries, fast fashion, its relation to the environment and unethical practices already constitutes a polarized topic. Therefore, although classical opinion dynamics models suggest that increased interaction rates would eventually lead to a consensus ([10], [14]), even on controversial issues, this has been challenged by vast empirical evidence of opinion polarization. We account for such developments by introducing a polarized agentset. This mechanism similar to existing ABM models on environment adjacent topics ([58]), where agents are influenced by others based on similarities in identity.
@@
The equation used for social interactions in a polarized society will be similar to that for a non-polarized society. However, the difference lies in the introduction of a tolerance threshold (denoted by $\tau\in[0.05,0.50]$), which represents how tolerant agents are toward opinions that deviate from theirs by more or less than the threshold. This is a further extension of the Friedkin–Johnsen framework. The tolerance threshold we introduce results in a opinion dynamic similar to that of the Bounded Confidence Model ([26]), where individuals are influenced only by others whose opinions differ by less than a specified threshold. We provide more details on the modified opinion-updating mechanism in the Appendix A.2. In summary, a lower threshold indicates less tolerance towards differing opinions. When an agent encounters an opinion that exceeds their tolerance threshold, the interaction becomes polarizing rather than unifying.
4.4.3 Social media influence.
@@
In real life, social media platforms adapt to each user, tailoring content based on their preexisting preferences and inclinations. Content one sees is content they already consume or are likely to consume. In our model, we implement such mechanisms by incorporating a feedback loop between the agent’s opinion and their social media platform. It is recognized that social media tends to increase polarization by creating “echo chambers” that limit exposure to information contradicting preexisting beliefs ([5]), and this is particularly the case for politicized topics ([37]). Moreover, social media can be biased towards pro-fast fashion or anti-fast fashion, which is something we consider in this model. Our choice to add a bias term rests on the fact that social media is found to both promote consumption ([20], [46]), but can also be a powerful tool that can be used to promote sustainable habits ([45], [55], [64]). As a result, we derive a function that takes an agent’s current opinion as an input an returns an “influence” opinion to the agent. This feedback loop is given by Eq. (3).
$SM\left(A_{t-1}\right)=b\cdot\left(A_{t-1}\right)^{3}+\frac{-3\cdot b}{2}\left(A_{t-1}\right)^{2}+\frac{3\cdot b}{4}\left(A_{t-1}\right)+\left(\frac{1}{2}-\frac{b}{8}\right)+\beta,$ (3)
were $b=50\cdot S_{sm}$, $S_{sm}$ stands for agent’s susceptibility towards social media, and $\beta\in[-0.30,0.30]$ is the bias. The rationale behind the choice of function for this feedback loop is detailed in Appendix A.2. The output varies for each agent’s opinion, reflecting the individual calibration of social media platforms. Once we have obtained the social media influence level from the function above, we can find the agent’s opinion after using social media, which is given by:
@@
where $A_{t-1}$ represents agent’s opinion at the previous time step, $A_{t}$ is the updated opinion, $S_{sm}$ stands for social media susceptibility of the agent, and $SM$ is the opinion that social media obtained from the feedback loop. The rationale behind this choice of function is detailed in the Appendix A.2.
4.4.4 Government interventions.
@@
$A_{t}=(1-S_{gov})\cdot g(A_{t-1},GOV)\cdot A_{t-1}+S_{gov}\cdot f(A_{t-1},GOV)\cdot GOV,$ (6)
where $GOV$ is the promoted opinion calculated with Eq. (5), $S_{gov}$ is agent’s susceptibility to government’s interventions, and $A_{t-1}$ is the opinion of an agent at time t1t-1. The functions $g$ and $f$ are exactly the same as described above for social media, and Eq. (6) is identical to Eq. (4) explained in Appendix A.2. Similar to social media, we aimed to prevent drastic changes in agents’ opinions and thus chose a more sophisticated function that allowed for a slower evolution of their viewpoints. Additionally, we implement a function that aims to mimic the diminishing impact of repeated exposure, a phenomenon known as “campaign fatigue” ([52]). This will affect agents’ susceptibility to government, $S_{gov}$, in a manner represented by Eq. (7).
@@
4.5 Structure overview and scheduling
Once agents and their friend groups are initialized, the simulation begins. It is set to run for 500 time steps, which was chosen as the limit because further iterations provided no qualitative differences in outcomes. Each time step, agents deterministically interact with some of their neighbors and social media once, and stochastically interact with government campaigns. It is important to note that these discrete time steps that do not necessarily correspond to real-world units such as days. This limitation reflects a common issue in ABMs known as temporal granularity, and the lack of longitudinal data prevents us from calibrating the model’s time scale. Consequently, the timing should be interpreted qualitatively rather than quantitatively. At each time step, we measure the agent’s probability to purchase fast fashion as the variable of interest, which does not necessarily correspond to the act of purchasing, but should rather be viewed as an individual’s acceptance of fast fashion and likelihood of purchasing it when faced with the decision. More information on the simulation can be found in Appendix A.3.
5 Model settings